# Contrastive Alignment with Semantic Gap-Aware Corrections in Text-Video Retrieval

### v4

@@ Abstract
Recent advances in text-video retrieval have been largely driven by contrastive learning frameworks. However, existing methods overlook a key source of optimization tension: the separation between text and video distributions in the representation space—referred to as the modality gap—and the prevalence of false negatives in batch sampling. These factors lead to conflicting gradients under the InfoNCE loss, impeding stable alignment. To mitigate this, we propose GARE—a Gap-Aware Retrieval framework that introduces a learnable, pair-specific increment $\Delta_{ij}$ between text $t_{i}$ and video $v_{j}$ to offload the tension from the global anchor representation. We first derive the ideal form of $\Delta_{ij}$ via a coupling multivariate first-order Taylor approximation of the InfoNCE loss under a trust-region constraint, revealing it as a key mechanism for resolving gradient conflicts by guiding updates along a locally optimal descent direction in the coupled optimization landscape. Due to the expensive cost of directly approximate $\Delta_{ij}$, we introduce a lightweight neural module conditioned on the semantic gap between each video-text pair, enabling structure-aware correction guided by gradient supervision. To further stabilize learning and promote interpretability, we regularize $\Delta$ via three components: a trust-region constraint regularization to prevent oscillations, a directional diversity term to expand the semantic difference space, and an information bottleneck over $\Delta$ to restrict redundant information. Experiments across four retrieval benchmarks show that GARE consistently improves alignment accuracy and robustness to noisy supervision, confirming the effectiveness of gap-aware tension unloading. Code is available at https://github.com/musicman217/GARE-text-video-retrieval.
School of Computer Science and Information Engineering,
Hefei University of Technology, Hefei, China
{j.xiao_hfut, chenghao}@mail.hfut.edu.cnzjsonghfut@gmail.comzdszds534@gmail.com{zzhu, lijia}@hfut.edu.cnhongrc.hfut@gmail.com
@@ 1 Introduction
Text-video retrieval (TVR) [48] is a fundamental task in video understanding, aiming to retrieve relevant videos given a text query [29, 28, 39, 15]. With the proliferation of short video platforms, this task has attracted growing research interest. In recent years, vision-language pretraining models such as CLIP [33] have shown great success in cross-modal representation alignment, demonstrating strong performance on various retrieval benchmarks. These models typically learn a shared embedding space by aligning visual and textual modalities through large-scale contrastive learning [43, 18, 8, 9, 17, 40, 19, 27], and have thus become a popular backbone in TVR systems.
Despite the empirical success of contrastive learning in text-video retrieval, existing methods largely overlook two fundamental sources of optimization tension. First, the significantly modality gap between text and video representations: the two modalities often reside in disjoint regions of the embedding space [27, 47], with markedly different semantic structures and high distributional divergence (e.g., large KL divergence between modality-wise feature distributions). This separation creates an inherent geometric and semantic mismatch that cannot be easily bridged by global metric learning. Second, the prevalence of false negatives: semantically similar but unlabeled text-video pairs—further injects noisy gradient signals during training [10, 7], as they are incorrectly treated as hard negatives. These two factors jointly induce unstable or conflicting optimization forces during contrastive learning, especially under the coupled softmax structure of InfoNCE [32]. Without an explicit mechanism to account for these challenges, models often converge to suboptimal solutions, particularly on semantically ambiguous or structurally complex samples.
@@
Figure 1: Tension and false negative challenge vs. our tension offloading strategy. (a) Due to modality gap [27] and false negatives [38], the gradients from negative samples partially overlap with the correct direction, leading to gradient tension around the anchor $t_{i}$. This restricts its optimization freedom and causes erroneous updates. (b) We offload part of the optimization pressure from $t_{i}$ to the learned increment $\Delta_{ij}$ by relaxing the gradient. This relieves gradient tension and reduces the impact of false negatives on $t_{i}$, as errors are partially absorbed by $\Delta$. Each $\Delta_{ij}$ captures a semantically meaningful correction of the text-video gap.
Figure 1(a) provides an overview of the optimization tension caused by modality gaps, where text and video embeddings lie in well-separated regions of the representation space—this repulsion tends to push $t_{i}$ away from the entire video manifold. As a result, the positive and negative terms impose a conflicting descent direction, both acting along the same region but with opposing effects. This constitutes a core source of optimization tension in contrastive learning. In addition, false negatives—semantically relevant but unlabeled pairs—introduce noisy repulsion that further distorts the optimization trajectory, especially in early training.
To address both issues, we introduce a pair-specific increment $\Delta_{ij}$ that acts as a learnable adjustment between $t_{i}$ and $v_{j}$. As shown in Figure 1(b), instead of enforcing all semantic correction on the anchor embedding $t_{i}$, we allow $\Delta_{ij}$ to absorb part of the optimization signal at the pair level. This design serves two complementary purposes: 1) it unloads the global optimization tension caused by the modality gap by decoupling the descent direction across different video-text pairs; 2) it buffers the local gradient noise introduced by false negatives, preventing anchor representations from drifting due to misaligned repulsion.
To ensure that $\Delta_{ij}$ contributes constructively to the loss landscape, we derive its ideal form via a first-order Taylor expansion of the contrastive loss under a trust-region constraint. While the exact solution is intractable due to inter-pair coupling, we approximate it using a lightweight neural module that conditions on the semantic gap $v_{j}-t_{i}$, allowing the model to proactively infer alignment adjustments before receiving backpropagated signals. This results in a closed-loop training dynamic: the model first estimates a gap-aware correction based on its prior belief, then updates this belief through feedback from contrastive supervision. Together, this enables more stable training and more accurate modeling of semantic structure under noise.
We further analyze the geometric behavior of the learned increment $\Delta_{ij}$. Empirically, we observe that $\Delta_{ij}$ tends to form an obtuse angle with the initial gap vector $v_{j}-t_{i}$, averaging around 1.6 radians. This indicates that the model is not merely copying the raw modality difference, but is instead learning a structure-aware correction that incorporates higher-level alignment priors. Moreover, we find that the Euclidean distance between paired embeddings increases after applying $\Delta_{ij}$, suggesting that the model leverages pair-level adjustment to release optimization tension and enable finer-grained control over semantic positioning on the unit sphere. Our contributions are as follows:
  • We analyze the gradient structure of the InfoNCE loss and reveal its inherent multi-variable coupling behavior via introduced multiple increments $\Delta$. By performing a first-order Taylor expansion under a local trust region, we derive an optimal update rule for each pairwise increment $\Delta_{ij}$ that aligns with the descent direction of InfoNCE.
  • Building on this formulation, we propose a gap-aware retrieval framework where each increment $\Delta_{ij}$ is predicted by a learnable network and directly participates in the forward pass. This design enables explicit offloading of gradient tension from text anchors, thereby mitigating misaligned updates caused by modality gap and false negatives.
  • We validate our method on four cross-modal retrieval benchmarks, i.e., MSR-VTT [45], DiDeMo [2], ActivityNet Captions [24], MSVD [6], demonstrating consistent improvements. Further analysis shows that the learned increments are semantically meaningful and geometrically structured, highlighting their role in stable and interpretable optimization.
2 Related work
Contrastive Learning and Modality Gap
Contrastive learning has become a foundational paradigm in multimodal representation learning. Wang and Isola [40] formalize contrastive learning via the principles of alignment and uniformity on the hypersphere, offering a geometric perspective on representation quality. Wang and Liu [38] show that contrastive loss is hardness-aware and temperature-sensitive, and reveal a trade-off between representation uniformity and semantic tolerance, highlighting the need to preserve meaningful structure among semantically similar samples. Liang et al.[27] investigate the modality gap in multimodal contrastive learning and attribute it to initialization imbalance and cone effects. False negatives have also been recognized as a key challenge in contrastive learning [10, 7], with solutions ranging from reweighting and elimination to dynamic detection and correction.
Text-Video Retrieval
Text-video retrieval is one of the prominent tasks  [13, 23, 37, 3, 16, 44, 31, 4, 21, 25, 41] in cross-modal learning. The majority of existing research  [26, 14, 22, 30, 42, 5, 46] in this area utilizes a mapping technique that aligns both text and video inputs within a shared latent space to facilitate direct similarity assessment. CLIP4Clip [29] is the first to adapt CLIP [33] for video-text retrieval via temporal frame aggregation. TS2-Net [28] improves temporal modeling through token shift and selection. X-Pool [15] introduces text-guided pooling to highlight salient video tokens. HBI [19] values possible correspondences between frames and words using Banzhaf interaction for sensitive and explainable cross-modal alignment. EMCL [19] introduces an expectation-maximization [11] framework to learn a compact latent space where video and text features are represented as linear combinations of shared bases. This decomposition reduces the rank of the latent space, mitigating the modality gap and enhancing semantic alignment. Unlike prior works that refine matching structures, our method analyzes the gradient form of InfoNCE and introduces a learnable gap-aware increment $\Delta_{ij}$ to offload optimization tension, enabling structured optimization in a trust-region-aware formulation.
@@ 3 Approach
We propose a gap-aware retrieval framework that introduces a learnable, pair-specific increment $\Delta_{ij}$ to adjust text features toward video embeddings or vice versa. To alleviate optimization tension and false negatives, we derive $\Delta_{ij}$ via a first-order Taylor expansion of the InfoNCE loss and approximate it using a neural module $\psi$ conditioned on the semantic gap. We further regularize $\Delta_{ij}$ in norm, direction, and distribution to enhance its structure and interpretability.
@@ 3.1 Preliminaries
Optimization Tension and First-order Increment Modeling
@@
$\nabla_{t_{i}}\mathcal{L}_{i}=\sum_{j=1}^{B}(p_{ij}-y_{ij})\cdot\left[\frac{v_{j}}{\|t_{i}\|\|v_{j}\|}-\cos(t_{i},v_{j})\cdot\frac{t_{i}}{\|t_{i}\|^{2}}\right],$ (3)
where $p_{ij}=\text{softmax}(\cos(t_{i},v_{j})/\tau)$ and $y_{ij}=\mathds{1}_{[j=i]}$. As observed in prior works such as [38], the magnitudes of the positive gradient coefficient $|p_{ii}-1|$ and the aggregate negative terms $\sum_{j\neq i}p_{ij}$ are comparable. In cases of strong modality gap, the attraction from the positive and the repulsion from negatives tend to pull in opposing directions, creating intrinsic gradient tension on $t_{i}$.
To alleviate both challenges mentioned above, we introduce a pair-specific increment $\Delta_{ij}$ that enables the model to locally adjust the relative positioning between each text-video pair. This increment serves a dual role: it offloads optimization tension by distributing alignment corrections across pairs, and buffers gradient noise by absorbing false-negative interference before it propagates into the anchor embedding. For clarity, we first consider the case where $\Delta_{ij}$ is applied to the text side, yielding an adjusted representation:
$t_{\Delta_{ij}}=t_{i}+\Delta_{ij}.$ (4)
This formulation allows for certain perturbations at the given anchor $t_{i}$ corresponding to different $v_{j}$ respectively. In practice, depending on dataset-specific modality characteristics, the same mechanism can also be applied to the video side (e.g., $v_{j}+\Delta_{ij}$). However, this flexibility raises a central question: how should each $\Delta_{ij}$ be optimized to effectively reduce the contrastive loss?
@@
To address this, we treat all $\Delta_{ij}$ as jointly optimized variables, and reinterpret the per-anchor contrastive loss $\mathcal{L}_{i}$ as multivariate function over the full set $\{\Delta_{ij}\}_{j=1}^{B}$:
$\mathcal{L}_{i}(\Delta_{i1},\dots,\Delta_{iB})=-\log\frac{e^{s_{ii}/\tau}}{\sum_{k=1}^{B}e^{s_{ik}/\tau}},\quad s_{ij}=\cos(t_{i}+\Delta_{ij},v_{j}).$ (5)
Unlike standard per-sample optimization, the softmax structure couples all $\Delta_{ij}$, meaning the gradient of any single increment depends on the values of the others. To capture this dependency, we perform a multivariate first-order Taylor expansion around a prior coupled state $\Delta_{i*}^{(0)}=\{\Delta_{ij}^{(0)}\}_{j=1}^{B}$, where all increments are nonzero. In practice, we treat this prior as the current state at iteration $t$ — i.e., we let $\Delta_{i*}^{(t)}:=\Delta_{i*}^{(0)}$ — and analyze the local descent behavior from this reference point. This converts our optimization into an iterative process, where each step refines the current set of coupled increments. The resulting linear approximation is:
@@
Crucially, this expansion preserves the softmax-induced coupling: each gradient term $\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ is evaluated assuming that all other increments $\Delta_{ik\neq j}$ remain fixed at their nonzero states. The resulting linearized loss defines a local descent landscape for each $\Delta_{ij}$. To ensure reliable updates within this approximation, we impose a trust-region constraint that bounds the update magnitude:
$\|\Delta_{ij}\|\leq\varepsilon,$ (7)
which reflects our belief that meaningful corrections should occur within a localized region around $t_{i}$. Under this constraint, the optimal update direction for each $\Delta_{ij}$ corresponds to steepest descent along its local gradient:
$\Delta_{ij}^{*(t+1)}=\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\cdot\frac{\nabla_{\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})}{\|\nabla_{\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})\|},$ (8)
where step size $\alpha_{ij}^{(t)}$ is analytically determined to ensure $\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon$ and derived via the Cauchy–Schwarz inequality to project onto the trust-region boundary (see Appendix D for details). This yields the steepest descent direction of the linearized loss, evaluated at a coupled batch state $\Delta_{i*}^{(t)}$, where all increments are fixed but nonzero. While conceptually similar to coordinate-wise proximal optimization, directly applying this solution is impractical due to two key challenges: 1) gradient coupling: each $\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ depends on all other $\Delta_{ik}$, making updates interdependent and unstable; 2) scale ambiguity: the optimal trust-region radius $\varepsilon$ may vary across pairs depending on semantic difficulty or training stage. To address this, we adopt a learnable network that directly predicts the coupled increment state $\Delta_{i*}^{(t)}$ based on the semantic gap $v_{j}-t_{i}$. The subsequent update to $\Delta_{i*}^{(t+1)}$ is implicitly performed via backpropagation on the InfoNCE loss. Since the gradient of the loss naturally aligns with the steepest descent direction, this formulation allows us to transform the iterative update rule into a trainable network prediction problem, enabling amortized, structure-aware optimization over the full batch of pairwise increments.

### v5

# Rebalancing Contrastive Alignment with Bottlenecked Semantic Increments in Text-Video Retrieval

@@ Abstract
Recent progress in text–video retrieval has been largely driven by contrastive learning. However, existing methods often overlook the effect of the modality gap, which causes anchor representations to undergo in-place optimization (i.e., optimization tension) that limits their alignment capacity. Moreover, noisy hard negatives further distort the semantics of anchors. To address these issues, we propose GARE, a Gap-Aware Retrieval framework that introduces a learnable, pair-specific increment $\Delta_{ij}$ between text $t_{i}$ and video $v_{j}$, redistributing gradients to relieve optimization tension and absorb noise. We derive $\Delta_{ij}$ via a multivariate first-order Taylor expansion of the InfoNCE loss under a trust-region constraint, showing that it guides updates along locally consistent descent directions. A lightweight neural module conditioned on the semantic gap couples increments across batches for structure-aware correction. Furthermore, we regularize $\Delta$ through a variational information bottleneck with relaxed compression, enhancing stability and semantic consistency. Experiments on four benchmarks demonstrate that GARE consistently improves alignment accuracy and robustness, validating the effectiveness of gap-aware tension mitigation. Code is available at https://github.com/musicman217/GARE-text-video-retrieval.
@@ 1 Introduction
Text-video retrieval (TVR) yu2017end is a fundamental task in video understanding, aiming to retrieve relevant videos given a text query luo2022clip4clip ; liu2022ts2 ; DRLTVR2022 ; gorti2022x . With the proliferation of short video platforms, this task has attracted growing research interest. In recent years, vision-language pretraining models such as CLIP radford2021learning have shown great success in cross-modal representation alignment, demonstrating strong performance on various retrieval benchmarks. These models typically learn a shared embedding space by aligning visual and textual modalities through large-scale contrastive learning wu2018unsupervised ; hjelm2018learning ; chen2020improved ; chen2021empirical ; he2020momentum ; wang2020understanding ; jin2022expectation ; liang2022mind , and have thus become a popular backbone in TVR systems.
Despite the empirical success of contrastive learning in text-video retrieval, two critical problems persist. First, the most challenge is optimization tension, arising from the modality gap: text and video embeddings typically occupy disjoint regions of the representation space liang2022mind ; yaras2024explaining , with markedly different semantic structures and high distributional divergence (e.g., large KL divergence between modality-wise feature distributions). As shown in Figure 1(a), this separation creates a structural conflict for a text anchor $t_{i}$: the gradient from its positive video $v_{i}$ attracts $t_{i}$ toward the video manifold, while gradients from all negatives $v_{j}$ repel it in the opposite direction, yielding nearly collinear but reversed forces. The second is the prevalence of false negatives: semantically similar yet unlabeled pairs are incorrectly treated as hard negatives chuang2020debiased ; chen2021incremental , introducing noisy gradients and aggravating misalignment. These two issues jointly limit the ceiling of semantic alignment, leading models to converge to suboptimal solutions.
@@
Figure 1: Tension and false-negative challenge vs. our offloading strategy. (a) Owing to the modality gap liang2022mind , gradients from negative samples overlap with the positive direction, creating optimization tension around the anchor $t_{i}$ and limiting its update freedom. (b) GARE offloads part of this optimization pressure from $t_{i}$ to learnable increments $\Delta_{ij}$, relaxing the gradient field and absorbing false-negative noise. Each $\Delta_{ij}$ encodes a semantically meaningful correction of the text–video gap.
Refer to captionRefer to caption
Figure 2: Mean and variance of summed gradient (top) and negative gradients (bottom) across 512 dimensions, showing collinear but opposite forces that largely cancel out.
We further analyze the optimization tension through aggregated gradient statistics across batches (Figure 2). In dimensions exhibiting significant gradient behavior, both positive and negative components reach values on the order of 40–60 (bottom of Figure 2), yet their sum—the actual gradient applied to $t_{i}$—remains close to 2–3 (top of Figure 2). This indicates that positive and negative signals not only oppose each other in direction but also nearly cancel in magnitude. As a result, text anchors undergo an in-place optimization behavior: their representations remain close to their initial positions throughout training, thereby limiting the attainable alignment performance of contrastive learning.
To address both issues, we introduce a pair-specific increment $\Delta_{ij}$ that acts as a learnable adjustment between $t_{i}$ and $v_{j}$. As shown in Figure 1(b), unlike the anchor embedding $t_{i}$, which aggregates gradients from all video pairs $(i,k)$, the increment $\Delta_{ij}$ only absorbs gradients transmitted from its corresponding pair $(i,j)$. This ensures that each $\Delta_{ij}$ captures a localized component of the optimization signal, while $t_{i}$ retains global supervision. As a result, gradients acting on $t_{i}$ are partially redistributed to $\Delta_{ij}$, effectively diluting the optimization tension and preventing anchors from being trapped in conflicting descent directions. Beyond relieving tension, this design also buffers the local gradient noise introduced by false negatives, since noisy repulsion from mislabeled pairs is absorbed at the pair level rather than directly perturbing anchor representations.
To guide $\Delta_{ij}$ toward constructive updates, we derive its local rule from a multivariate first-order Taylor expansion of the contrastive loss under an $\ell_{2}$ trust-region constraint. This linearization defines a descent space in which, once a constraint radius is specified, the steepest-descent direction at the coupled state becomes unique and preserves the local relative-ranking structure of InfoNCE. We implement this process as a pair-specific, amortized update using a lightweight module $\psi$ conditioned on the semantic gap $(v_{j}-t_{i})$. The module is optimized through backpropagation across batches, while a norm-based prior regularizes the magnitude of $\Delta$, serving as an implicit trust-region radius regularization. To ensure semantically stable and generalizable updates, we formulate the learning of $\Delta_{ij}$ as a deterministic variational information bottleneck: the module $\psi$ outputs $\Delta_{ij}$, and a Jensen-relaxed KL bottleneck term balances informativeness and compression.
On the MSR-VTT 1k-A validation set, we observe that the overall cosine similarity of positive pairs decreases relative to the baseline. This indicates that releasing optimization tension via $\Delta$ encourages samples to spread with larger angular separations, thereby improving uniformity on the unit hypersphere wang2020understanding . In addition, we find that the Euclidean distance between paired embeddings increases after applying $\Delta_{ij}$, suggesting that pair-level adjustments help alleviate tension and provide finer control over semantic positioning.
Our contributions are threefold: 1) We analyze the gradient structure of InfoNCE and reveal its inherent multi-variable coupling by introducing pairwise increments $\Delta_{ij}$. A multivariate first-order Taylor expansion within a trust region yields a update rule for each $\Delta_{ij}$ consistent with the InfoNCE descent direction. 2) We propose a Gap-Aware Retrieval (GARE) framework, where a learnable network predicts pair-specific increments $\Delta_{ij}$ and integrates them into the forward pass to offload optimization tension while mitigating noise from false negatives. We also introduce a relaxed variational information bottleneck (VIB) objective that regularizes $\Delta_{ij}$, balancing informativeness and compression. 3) Experiments on four text–video retrieval benchmarks, i.e., MSR-VTT xu2016msr , DiDeMo anne2017localizing , ActivityNet Captions krishna2017dense , and MSVD chen2011collecting , showing consistent improvements, and further analyses confirm that the learned increments are semantically meaningful and geometrically structured.
@@ 2.1 Preliminaries
Optimization Tension and First-order Increment Modeling.
@@
$\nabla_{t_{i}}\mathcal{L}_{i}=\sum_{j=1}^{B}\underbrace{\frac{(p_{ij}-y_{ij})}{\tau}}_{\text{weight}\ \frac{\partial\mathcal{L}_{i}}{\partial s_{ij}}\in\mathbb{R}}\cdot\underbrace{\left[\frac{v_{j}}{\|t_{i}\|\|v_{j}\|}-\cos(t_{i},v_{j})\cdot\frac{t_{i}}{\|t_{i}\|^{2}}\right]}_{\text{gradient basis }\ \frac{\partial s_{ij}}{\partial t_{i}}\in\mathbb{R}^{D}},$ (3)
where $p_{ij}=\text{softmax}(\cos(t_{i},v_{j})/\tau)$ and $y_{ij}=\mathds{1}_{[j=i]}$. Here, the weight term $\frac{\partial\mathcal{L}_{i}}{\partial s_{ij}}\in\mathbb{R}$ controls the relative strength of each pair, while the gradient basis $\frac{\partial s_{ij}}{\partial t_{i}}\in\mathbb{R}^{D}$ specifies the update direction. Under a strong modality gap, the positive and negative gradient bases are nearly collinear but reversed, and the positive weight $|p_{ii}-1|$ is of comparable magnitude to the aggregate negative weight $\sum_{j\neq i}p_{ij}$ wang2021understanding . Together, these conditions cause the gradients to largely cancel along nearly the same axis, leaving only a small residual update for $t_{i}$.
To alleviate both challenges, we introduce a pair-specific increment $\Delta_{ij}$ that locally adjusts the relative positioning of each text-video pair. This increment serves two purposes: it offloads optimization tension by redistributing part of the gradients originally acting on $t_{i}$ to the pair level, and buffers false-negative noise before it propagates to the anchor embedding. For clarity, we first consider applying $\Delta_{ij}$ to the text side, yielding an adjusted representation $t_{\Delta_{ij}}=t_{i}+\Delta_{ij}$. The same mechanism can also be applied to the video side (e.g., $v_{\Delta_{ij}}=v_{j}+\Delta_{ij}$), depending on dataset characteristics. This flexibility raises a central question: how should each $\Delta_{ij}$ be optimized to effectively reduce the contrastive loss?
@@
To this end, we treat all $\Delta_{ij}$ as jointly optimized variables, and reinterpret the per-anchor contrastive loss $\mathcal{L}_{i}$ as multivariate function over the full set $\{\Delta_{ij}\}_{j=1}^{B}$:
$\mathcal{L}_{i}(\Delta_{i1},\dots,\Delta_{iB})=-\log\frac{e^{s_{ii}/\tau}}{\sum_{j=1}^{B}e^{s_{ij}/\tau}},\quad s_{ij}=\cos(t_{i}+\Delta_{ij},v_{j}).$ (4)
Unlike standard per-sample optimization, the softmax structure couples all $\Delta_{ij}$, meaning the gradient of any single increment depends on the values of the others. To capture this dependency, we perform a multivariate first-order Taylor expansion around a prior coupled state $\Delta_{i*}^{(0)}=\{\Delta_{ij}^{(0)}\}_{j=1}^{B}$, where all increments are nonzero. In practice, we treat this prior as the current state at iteration $t$ (i.e., we let $\Delta_{i*}^{(t)}:=\Delta_{i*}^{(0)}$) and analyze the local descent behavior from this reference point. This converts our optimization into an iterative process, where each step refines the current set of coupled increments. The resulting linear approximation is:
@@
Crucially, this expansion preserves the softmax-induced coupling: each gradient term $\nabla_{\Delta_{ij}}\mathcal{L}_{i}$ is evaluated assuming that all other increments $\Delta_{ik\neq j}$ remain fixed at their nonzero states. The resulting linearized loss defines a local descent landscape for each $\Delta_{ij}$. To ensure reliable updates within this approximation, we impose a trust-region constraint that bounds the update magnitude: $\|\Delta_{ij}\|\leq\varepsilon$, which reflects our belief that meaningful corrections should occur within a localized region around $t_{i}$. Under this constraint, the optimal update direction for each $\Delta_{ij}$ corresponds to steepest descent along its local gradient:
$\Delta_{ij}^{(t+1)}=\Delta_{ij}^{(t)}-\alpha_{ij}^{(t)}\cdot\frac{\nabla_{\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})}{\|\nabla_{\Delta_{ij}}\mathcal{L}_{i}(\Delta_{i*}^{(t)})\|},$ (6)
where step size $\alpha_{ij}^{(t)}$ is analytically determined to ensure $\|\Delta_{ij}^{(t+1)}\|\leq\varepsilon$ and derived via the Cauchy–Schwarz inequality to project onto the trust-region boundary (see Appendix H for details). This yields the steepest descent direction of the linearized loss, evaluated at a coupled batch state $\Delta_{i*}^{(t)}$, where all increments are fixed but nonzero. However, this update rule has two main limitations: (i) it only guarantees descent within the current batch, lacking cross-batch generalization; and (ii) it suffers from scale ambiguity, as the optimal trust-region radius $\varepsilon$ should vary across pairs with semantic difficulty and training stage. To overcome these issues, we adopt a learnable network $\psi$ that directly predicts the coupled increment state $\Delta_{i*}^{(t)}$ from the semantic gap $v_{j}-t_{i}$ (or $t_{i}-v_{j}$), thereby establishing cross-batch coupling among increments. The subsequent update to $\Delta_{i*}^{(t+1)}$ is implicitly performed via backpropagation on the InfoNCE loss. Since the loss gradient naturally aligns with the steepest descent direction, this formulation amortizes the iterative update process into a trainable prediction problem, enabling structure-aware and generalizable optimization over the full set of pairwise increments.